### v1

# Augmenting the action space with conventions to improve multi-agent cooperation in Hanabi

Abstract
The card game Hanabi is considered a strong medium for the testing and development of multi-agent reinforcement learning (MARL) algorithms, due to its cooperative nature, hidden information, limited communication and remarkable complexity. Previous research efforts have explored the capabilities of MARL algorithms within Hanabi, focusing largely on advanced architecture design and algorithmic manipulations to achieve state-of-the-art performance for a various number of cooperators. However, this often leads to complex solution strategies with high computational cost and requiring large amounts of training data. For humans to solve the Hanabi game effectively, they require the use of conventions, which often allows for a means to implicitly convey ideas or knowledge based on a predefined, and mutually agreed upon, set of “rules”. Multi-agent problems containing partial observability, especially when limited communication is present, can benefit greatly from the use of implicit knowledge sharing. In this paper, we propose a novel approach to augmenting the action space using conventions, which act as special cooperative actions that span over multiple time steps and multiple agents, requiring agents to actively opt in for it to reach fruition. These conventions are based on existing human conventions, and result in a significant improvement on the performance of existing techniques for self-play and cross-play across a various number of cooperators within Hanabi.
keywords:
Multi-agent Reinforcement Learning, Deep Learning, Hanabi, Agent Cooperation, Cross-play
1 Introduction
Reinforcement learning (RL) holds promising potential to address a large variety of problems where artificial agent operation offers a significant improvement over alternative methods, such as hard-coded algorithms or solutions [1]. For certain real-world problems, single-agent operation is not optimal (or even possible) and the incorporation of multiple agents would be beneficial (or necessary). Scenarios such as autonomous vehicles navigating terrain [2], guided drone swarms [3], mapping verbal instructions to executable actions [4], or the switching of railway lines benefit from the use of multiple agents. RL can help agents to effectively cooperate within these multi-agent scenarios by learning from past and or simulated experiences.
Unfortunately, the introduction of multiple agents into an environment typically increases the complexity of the problem exponentially. It introduces a moving target learning problem [5], since all the agents must learn simultaneously. The reason for this is that each individual agent’s policy changes over time, which in turn causes a non-stationary environment. This often inhibits all the agents from developing effective policies and can lead to undesired behaviour [5].
One of the most straightforward solutions to multi-agent reinforcement learning (MARL) is the combination of deep Q-networks (DQNs) and independent Q-learning [6], where each agent independently and simultaneously learns its own action-value function (or Q-values [^1] Action values ($Q$) refer to the value assigned to a certain action $a$ within a given state $s$.) while interacting with the same environment. However, this strategy does not fair well when paired with partial observability [^2] Partial observability refers to problems where the full state space is hidden from each individual agent, resulting in each agent having their own unique perspective of the problem. [7, 8]. Hausknecht and Stone [9] have shown that recurrent neural networks offer an improved solution to MARL problems containing partial observability. These networks incorporate a built-in short term memory over which experiences are unrolled (or combined) to form longer sequences. This is often combined with deep Q-learning’s feed forward neural network to produce deep recurrent Q-learning [9].
Various research efforts follow a similar path as Hausknecht and Stone, focusing on architectural advancements through the means of neural network layer manipulations, and complex algorithms to improve the estimation and assignment of action-values or policies [10, 11, 12]. This often leads to convoluted solution strategies that are difficult to implement and computationally expensive [13, 14]. An alternative solution strategy would be to reconsider the problem dynamics and discover ways of incorporating existing domain knowledge into RL algorithms. Existing domain knowledge can be incorporated into RL and MARL using reward shaping [15, 16, 17], which focuses on manipulating the reward signal to encourage certain behaviour. Alternatively, agents can use state-space augmentations, such as auxiliary tasks [18], to imply or extract additional information (or features) of the problem setting. Sutton et al. [19] have shown that options offer an additional strategy for incorporating domain knowledge into RL by changing the action space of an agent. Options allow an agent to solve problems on a higher level by extending the action space to include advanced temporal actions that range over multiple time steps (also referred to as macro-actions[19].
Games are often used as a medium for testing and evaluating RL algorithms, since they incorporate real world problems with well-defined rules and a clear metric for measuring performance, where examples include Go [20], Backgammon [21], and Dota 2 [22]. Hanabi is a cooperative card game containing partial observability and limited communication, requiring players to logically reason over the intentions and actions of their cooperators, a concept known as theory of mind [23]. Human players require the development of conventions [^3] Philosopher David Lewis defines conventions as an arbitrary, self-perpetuating solution to a recurring coordination problem [24]. in order to solve the problem effectively. Conventions have evolved with humanity throughout the years and range from driving on a certain side of the road to social conventions or norms [24].
Most research efforts on the Hanabi problem use complex architecture design with advanced algorithms, often focusing on estimating and calculating the beliefs of other agents [25, 13]. However, one avenue of the Hanabi problem yet to be explored is that of conventions, and how to incorporate existing human conventions into MARL solution strategies. In this paper, we propose a method to incorporate human conventions into MARL through the use of artificial conventions, which act as cooperative actions that span over multiple time steps and agents, and show how it significantly improves on the performance of MARL agents within Hanabi for self-play [26] as well as cross-play [27] scenarios. Our approach shares similarities with options, due to the multi-time step extension of actions, however it is fundamentally built on a different concept.
1.1 Related Work
A popular solution strategy to RL problems introduced by Hessel et al. [28], called Rainbow, combines various advancements made to deep Q-learning and has proven to offer significant performance gain when faced with large discrete action spaces [29, 30, 31]. A natural extension of Rainbow to multi-agent systems, is the implementation of independent Rainbow agents, referred to as multi-agent Rainbow (MA-Rainbow) [32]. However, Rainbow introduces a plethora of new hyperparameters which can often result in suboptimal policy development for cooperative settings [33], and is known for falling into suboptimal local minima [34, 35, 36].
Hanabi was first proposed as a viable medium for MARL by Bard et al. [36], with focus placed on the philosophical ideas and challenges found within the problem setting and how they translate to real world scenarios. Bard et al. conducted initial tests for three different MARL algorithms, namely MA-Rainbow, actor-critic-Hanabi-agent (ACHA) [36], and Bayesian action decoder (BAD) [25], and compared them to state-of-the-art handcrafted bots, such as HatBot [37] and WTFWThat [38]. The MARL solutions achieved an acceptable score with the best performing agent achieving 16.8/25 for the most difficult scenario of five players, but still significantly falls behind the best handcrafted bot which achieves a score of 24.89/25.
The simplified action decoder (SAD), introduced by Hu and Foerster [13], aims to solve a similar goal as the BAD algorithm, but with the main benefit of reducing the effect exploration actions have on the public beliefs of agents [13]. They improve on the performance of the BAD algorithm, while also extending these concepts to more than just two players. Recurrent neural networks [9] are used to track the public beliefs and an auxiliary task [18] is trained to predict if a card is playable, discardable or unknown. Hu and Foerster found that these auxiliary tasks only benefited the two player scenario while drastically hurting the 3–5 player performance [13]. They are able to achieve good performance for 2–5 players in Hanabi, and close the gap between RL agents and handcrafted solutions by achieving a score of 22.06/25 in five player Hanabi. However, the SAD algorithm is computationally expensive, requiring a minimum of 40 CPU cores, 2 high-end GPUs and 256 gigabytes of RAM, while still requiring billions of samples and a wall time of 72 hours [13].
Lerer et al. [14] further builds on the concept of the SAD algorithm by introducing multi-agent tree search (MATS) [39], and manage to achieve state-of-the-art performance for 2–5 player Hanabi. The agents manage to beat the current best handcrafted bots in the two player scenario, with the handcrafted bots only outperforming the agents in five player by 4%. However, just like its predecessor, the SAD-MATS algorithm is computationally expensive and requires a substantial amount of training data and wall time to achieve these results [14].
Hu et al. [27] also builds on the concept of the SAD algorithm, but rather than achieving state-of-the-art performance for self-play agents, they focus on the agent’s ability to cooperate with a variety of partners–a concept known as cross-play. This is achieved using other-play, an architecture for multi-agent systems specifically designed to break the ties formed between self-play agents, allowing them to produce more robust strategies. This allows the agents to cooperate with agents from different training runs (or regimes), and even human cooperators [27]. Even though the other-play agents are able to achieve effective cross-play performance when paired with a variety of agents in a two player scenario, the research does not focus on higher player counts where more advanced cooperation strategies are required. Additionally, other-play builds on the existing architecture of the SAD algorithm leading to increased computational complexity and cost [27].
In contrast to value-based methods, Yu et al. [40] show that Multi-agent Proximal Policy Optimization (MAPPO) offers similar performance to value-based alternatives in four different cooperative tasks, namely multi-agent particle world environment (MPE) [41], StarCraft micromanagement challenge (SMAC) [42], Google research football (GRF) [43], and the Hanabi challenge [36]. MAPPO is able to match, and in some instances outperform the capabilities of leading edge value-based methods, such as Qplex [44], RODE [45], and value decomposition networks (VDNs) [10]. It shows promising results in Hanabi and often outperformed MA-Rainbow, but is sample inefficient requiring a substantial amount of training steps (more than 10 billion) to achieve good performance.
1.2 Summary of Contributions
Rather than focusing on complex architecture design, this research shifts focus to incorporating existing domain knowledge of the Hanabi problem into MARL. Hanabi has an active and dedicated community with players constantly searching for new conventions and ways to reliably beat the game. These conventions are often based on principles, and have generally been standardised within the Hanabi community. Our main contribution is to show how these existing human conventions can be implemented in a MARL scenario using artificial conventions, which act as a special form of cooperative actions and action space augmentations.
One of the key insights to this approach is our “subscribing” technique for convention continuation, where an agent can initiate a convention (based on a certain human convention) and subsequently a new action appears to the other agents to “subscribe” or continue with this convention. This strategy requires that agents opt in to a certain convention in order for it to reach completion, and allows an agent to halt a certain inferior convention, and/or initiate an improved convention, based on their unique observation. To this end, each convention contains an initial condition acting as the start of a convention, subsequent continuation conditions for other agents to opt in, a termination condition for a terminating action, and finally a policy that translates each aspect of the convention into an environment action based on the current agent’s observation.
Our approach shares similarities with that of multi-agent options, but addresses a problem where agents must often wait for other agents to complete their options before they can initiate a new option (which can include a cooperative option) [46]. Fundamentally, our approach is built on a different concept, due to the cooperative requirement of conventions to reach fruition, and their ability to convey ideas or intentions through implicit communication based on mutually agreed upon “rules”. Therefore, it is important to note that our approach does not introduce additional explicit communication, or additional communication channels for that matter, but rather allows for implicit communication through the nature of conventions.
Conventions result in a large performance uplift for existing MARL techniques, in our case multi-agent Rainbow, by significantly decreasing the training time and, in the case of 3–5 player Hanabi, increasing the converged performance. Additionally, since the agents are learning from a communal and well-defined list of conventions, the agents are not just able to cooperate in self-play, but also in a cross-play setting with agents from different training runs (or regimes). Cross-play has been shown to be an important research area for MARL since it opens the door to agents being able to cooperate with never-before-seen teammates, whether they are agents from other regimes, agents with different architectures, or humans [47, 48].
1.3 Paper Outline
In Section 2 we introduce the background and theory of RL, MARL, and options. In Section 3, we discuss the Hanabi environment as well as existing solution strategies, with focus placed on human conventions, in order to give context for our approach. This is followed by the discussion on conventions and action space augmentation with conventions in Section 4. Finally, we present and discuss our results for self-play and cross-play for all player counts in Hanabi, in Section 5.
2 Multi-agent Reinforcement Learning and Options
Reinforcement learning aims at solving decision-making problems and is built on the concept of Markov decision processes (MDPs) [49]. Multi-agent reinforcement learning introduces more than one agent into the environment and often incorporates partial observability, and thus is built on the concept of decentralised partially observable Markov decision processes (Dec-POMDPs) [50]. The Dec-POMDP framework consists of
$\langle S,A,T,O,P,R,n,\gamma\rangle,$ (1)
where $S$ denotes the global state space and $A$ represents the joint-action space of the $n$ agents at time $t$. The observation space $O$ consists of the local observation of each agent $i$ at time $t$ within the global state $S$ ($O^{i}_{t}=O(S;i;t)$). After $A$ is sent to the environment, the global state transitions from $S$ to $S^{\prime}$ given the state transition probability $T$, and similarly $P$ represents the observation transition probability to transition from $O$ to $O^{\prime}$. This results in the environment producing $R$, which donates the immediate reward for each agent action that contributed to a global state change at $t$, with $\gamma$ denoting the discount factor for future reward. These rewards are either combined with equal weighting (summed), or can have their own discount factor $\gamma_{r}$, and is referred to as the forward accumulated reward [50].
In value-based RL we can use (1) to calculate the value functions ($Q$) which acts as a quantitative measure for the desirability of a state. In tabular methods these values are represented using a lookup-table, while deep RL methods use powerful function approximators to estimate these values.
2.1 Independent Q-learning
Q-learning is an off-policy, temporal difference (TD) control algorithm that learns the action-value function $Q(S,A)$ by directly approximating the optimal action-value function $Q^{*}(S,A)$, independent of the policy being followed [49]. The update step for the action-value function is defined as
$Q(S_{t},A_{t})\leftarrow{}$ $Q(S_{t},A_{t})+\alpha\bigr{[}R_{t+1}+\gamma\max_{a}Q(S_{t+1},a)-Q(S_{t},A_{t})\bigr{]},$ (2)
where $\alpha$ is the learning rate and $\gamma$ the discount factor of future rewards [49]. This can be extended to MARL with independent Q-learning, where each agent has their own action-value function and uses their individual experiences in their update steps [51]. The pseudocode for independent Q-learning in a turn-based scenario is shown in Algorithm 1, where each agent $i$ uses their own observations $O^{i}_{t}$ to update their action-value functions $Q^{i}$. Note that due to the unique nature of turn-based settings, a special type of forward accumulated reward (FAR) must be used which consists of the one round return resulting from each agent’s action, i.e. $\sum_{k=0}^{P-1}R_{t+1+k}$.
Algorithm 1 Independent Q-learning in a turn-based environment
[…]
2:Initialise all $Q(O,A)\leftarrow 0$
3:for each episode do
4:     Reset environment and set total time $t\leftarrow 0$
5:     while $S_{t+1}$ is not terminal do
6:         for all players $i$ from 00 to $P-1$ do
7:              $A^{i}_{t}=\text{argmax}_{a}Q^{i}(O^{i}_{t},a)$triangleright Choose $A^{i}_{t}$ using $O^{i}_{t}$ and the policy derived from $Q^{i}$
8:              Take action $A^{i}_{t}$ in the environment
9:              Receive and store $R_{t+1}$ and $O^{i}_{t+1}$, with $O^{i}_{t+1}\in S_{t+1}$, from environment
10:              $\sum_{k=0}^{P-1}R_{t+1+k}$triangleright Calculate the FAR based on the rewards for the next round of actions resulting from $A^{i}_{t}$
11:              $Q^{i}(O^{i}_{t},A^{i}_{t})\leftarrow Q^{i}(O^{i}_{t},A^{i}_{t})+\alpha\bigr{[}\sum_{k=0}^{P-1}R_{t+1+k}+\gamma\max_{a}Q^{i}(O^{i}_{t+P},a)-Q^{i}(O^{i}_{t},A^{i}_{t})\bigr{]}$
12:              $t\leftarrow t+1$
13:         end for
14:     end while
15:end for
Instead of distinct policies, independent Q-learning agents can also make use of a shared policy, especially when the environment contains symmetries [51]. This allows the update step (2) to be updated more frequently by using the experiences of all the independent agents. This has proven to offer significant performance gain and allow agents to learn more effectively [51].
2.2 Deep Q-learning
Deep Q-learning is an extension of tabular Q-learning where artificial neural networks (referred to as deep Q-networks or DQNs) are used as non-linear function approximators [52]. Deep Q-learning implements an experience replay memory to store the experiences $e=\langle S_{t},A_{t},R_{t+1},S_{t+1}\rangle$ [53], which in turn is sampled in random batches $b$ to remove correlation within the observation sequence, and thereby smoothing over the changes within the data distribution [52].
Deep Q-learning usually incorporates a policy network and a target network [52]. The policy network is used to select actions and utilises the update step, while the target network serves as a baseline when calculating the TD error. The target network is updated periodically with the policy network to reduce correlation with the target. The weights of the policy network are updated using backpropagation with the goal of minimizing the TD error, which is defined as
$L_{j}(\theta_{j})={}$ $\mathbb{E}_{b}[(R_{t+1}+\gamma\max_{a}Q(S_{t+1},a;\theta_{j}^{-})-Q(S_{t},A_{t};\theta_{j}))^{2}],$ (3)
where $\theta_{j}$ and $\theta_{j}^{-}$ are the weights of the policy-and target network, respectively, at iteration $j$ [52]. In practice the backpropagation and calculation of the TD error is usually handled by an optimizer, such as the Adam optimizer [54]. Similar to tabular Q-learning, deep Q-learning can be extended to MARL using independent deep Q-learning [6], and can also make use of a shared policy.
2.3 Options
Options are built on the concept of semi-MDPs, where actions can have varying lengths over multiple time steps $t$ as the environment transitions from $S$ to $S^{\prime}$ [19]. When an agent choses an option $\omega$, the option executes stochastically according to a policy $\pi_{\omega}$. This policy allows for existing domain knowledge to be incorporated into RL techniques. For example, in the case of a robot navigating terrain, a potential option can include: {move around boulder}, and once the agent choses that option, an object avoidance controller takes over executing policy $\pi_{\omega}$.
Once policy $\pi_{\omega}$ has concluded, the agent can choose a new option, effectively simplifying the problem at hand by solving it on a higher level, since the agent no longer has to focus on learning the primitive actions [^4] Primitive actions refer to the raw actions of the environment, for example: move up, move left, move right, move down, etc. as well as the overarching problem simultaneously. Thus, an option $\omega$ is defined as
$\langle I_{\omega},\pi_{\omega},\beta_{\omega}\rangle,$ (4)
where $I_{\omega}$ is the initial condition that must be met for an option to become available in a certain state ($S_{t}\in I_{\omega}$), $\pi_{\omega}$ the policy according to which the option executes, and $\beta_{\omega}$ the termination condition to which the option is executed [19]. The policy $\pi_{\omega}$ can also be defined using RL, a process known as intra-option learning [19]. This has further lead to the development of Option Discovery, which uses a Laplacian framework to allow an agent to define and learn its own options as well as the intra-option policies while interacting with the environment [55]. Options have also been extended to MARL with Amato et al. [46] introducing their macro-action Dec-POMDP (MacDec-POMDP) and showing how options affect the Dec-POMDP structure.
Recently, Chen et al. [56] explored option discovery within MARL and introduced their own variation on the Laplacian framework using Kronecker graphs. Options within MARL poses unique challenges due to the temporal disconnect of the actions as a direct result of them having varying lengths. The actions are also chosen asynchronously, which further adds to the complexity and non-stationarity of the environment [46]. Additionally, agents in a synchronous setting must often wait for other agents to complete their options before they can initiate a new option, especially if that new option is a cooperative one [46].
3 Hanabi
The Hanabi problem along with how human conventions are used to solve the problem effectively, must first be discussed in order to give context for our novel approach. Hanabi is a 2–5 player card game, best described as a cooperative solitaire [36]. Players have a hand of five cards for two and three player, and four cards for four and five player. Each card has a suit/colour (red, yellow, green, white, or blue) and a rank (1 to 5). The deck comprises 50 cards total, with 10 cards for each suit with a distribution of three 1’s, two 2’s, 3’s, and 4’s, and finally only one 5. The aim of the game is to stack these suits in numerical order from 1 to 5, however players cannot see their own cards, only the cards of their fellow players.
The players take consecutive turns, and on each turn a player can either play, discard, or hint. Hinting involves revealing all the cards in another player’s hand matching a certain rank or a certain suit, and consumes one of the limited (and shared) hint tokens. In the centre is a stack for each suit where the players must play their cards, and a successful play entails playing a card that follows the current card on top of a stack (starting at 0). If the card played does not follow the current card on a stack, the play was unsuccessful (called a misplay) and the players lose one of their shared life tokens. Discarding involves removing a card from the current player’s hand and adding it to the discard pile, effectively removing it from the game, while also replenishing a hint token. An example game is shown in Fig. 1.
Figure 1: An example game of Hanabi as seen from the perspective of player 1. It is now player 1’s turn, and they can take the hint, play or discard action.
At the start of the game the players have a shared total of three life tokens and eight hint tokens. If all the hint tokens are depleted, a player cannot take the hint action and must either play or discard a card from their hand. The players are awarded a shared score depending on the number of cards successfully placed on each stack. If the players manage to build the stacks to the maximum of 5 each, the game ends and the players receive a perfect score of 25/25. Alternatively, if the players lose all of their life tokens (called “bombing out”), or if the deck has been depleted, the game also ends. Note, bombing out results in a score of 0/25 independent of how many cards were played successfully. A discard action, or a misplay, will result in the card being removed from the game, and that player must draw a new card from the deck to replenish the missing card (which is also hidden from current the player).
Due to the partial observability imposed by a player’s hand being hidden, as well as the limited communication channel in the form of hinting, Hanabi offers a very interesting and useful challenge for MARL agents to solve [36]. Actions are highly correlated and players must coordinate effectively to achieve success, often requiring advanced strategies and reasoning over the intentions of other player’s actions to reliably beat the game.
3.1 Human Conventions in Hanabi
Due to the nature of Hanabi, players have too few hint tokens in order to effectively convey all the needed information (colour and rank) of the 25 playable cards [36]. This becomes even more challenging as the number of players increase and hint tokens become more valuable, as well as the number of playable turns decreasing [^5] As the number of players increase, the cards remaining in the deck decrease, and a single player has less playable turns.. Hints must therefore be used wisely, and must convey information about more than just one card. This is possible since more than one card with a similar rank or suit can be touched (or revealed) by a single hint, along with negative information [^6] Negative information refers to implied information about non-focus cards, e.g. if a player hints that two cards are green, the other cards must therefore be non-green. being given about all the other cards in a hand. Unfortunately, this is not enough, and players who don’t use additional strategies will struggle to reliably achieve a score above 15/25.
Players, therefore, require a way to convey additional information through implicit communication, while still remaining within the rules of the game. As stated by Bard et al. [36]: “This implicit information is not conveyed through the impact that an action has on the environment (i.e., what happens) but through the very fact that another player decided to take this action (i.e., why it happened)”. Players can then reason over the actions of others, and implicitly communicate through them. This is done through the use of conventions, or mutually agreed upon “rules”, that players develop outside the game. For example, in Fig. 1 the players have a convention where an ambiguous hint has the focus placed on their left-most card, therefore player 1 can hint to player 2 that they have two red cards in position one and three respectively, and player 2 will know (as a result of their convention) that the left-most card must be the playable red 1. Since the game’s release, players have developed their own intricate and extensive conventions, often specific to their group of friends or cooperators, examples include the Board Game Arena [^7] https://forum.boardgamearena.com/viewtopic.php?t=5252 and the H-Group [^8] https://hanabi.github.io/, with the H-Group being the most widely used and adapted.
4 Augmenting the Action Space with Conventions
Existing MARL approaches specific to Hanabi naturally develop conventions, but these conventions are nonsensical and vary greatly from run to run [27, 47, 48]. We propose incorporating existing human conventions into MARL algorithms using artificial conventions, which act as advanced cooperative actions that span over multiple time steps and agents. Conventions require that all the agents involved participate or “subscribe” in order for it to reach fruition, however the agents cannot communicate directly which convention is being started or followed.
Conventions can be divided into two categories, namely available conventions and active conventions. We define a convention $C$ as
$C=\langle I,\pi_{I},M,\pi_{M},F,\pi_{F}\rangle,$ (5)
where the initial condition $I$ must be met for a convention to become available ($O_{t}\in I$), and if an agent chooses to initiate $C$, the policy $\pi_{I}$ determines the corresponding environment action to start the convention. The convention will then become active, and another agent can “subscribe” to $C$ if they meet the condition $M$, with $\pi_{M}$ determining their environment action to continue $C$. Eventually an agent will have the opportunity to perform a unique continuation action to complete $C$, but only if they meet the final condition $F$, with $\pi_{F}$ determining their environment action to terminate $C$ and allow it to reach completion. For the sake of simplifying this discussion, and to be consistent with our implemented Hanabi conventions in later sections, we will focus on two-step conventions, where one agent initiates a convention $C$ and another agent must opt in to complete the convention, i.e. $C=\langle I,\pi_{I},F,\pi_{F}\rangle$. The concept behind conventions is best described at the hand of an example:
Conventions Example. In Fig. 1 each player is controlled by an agent, with the current player (referred to as the active player) being agent 1. The convention at hand, which we will refer to as $c_{1}$, states that if an ambiguous hint is given, the focus of that hint is the left-most (or newest) card [^9] This is a common human convention and is based on the ”Single card focus” principle as discussed by the H-Group.. Since the condition $I_{c1}$ is met, i.e. player 2 has two cards in their hand that share the same colour or rank and the left-most one is playable, the action {start $c_{1}$} is available for agent 1. Agent 1 chooses to initiate $c_{1}$ which in turn gets translated by the policy $\pi_{Ic1}$ to an environment action of {hint red to player 2}, ending player 1’s turn. It is now player 2’s turn, and they have just received an ambiguous colour hint from player 1, and given their current observation they can derive that $c_{1}$ is currently active. This triggers the subscribing (and also completing) condition $F_{c1}$, where agent 2 can now opt in to continue (and complete) $c_{1}$. When agent 2 chooses to subscribe to $c_{1}$, their action is translated by the policy $\pi_{Fc1}$ to an environment action of {play card in position 1}, subsequently ending player 2’s turn and completing $c_{1}$.
It is important to note that conventions do not have to be sequential, e.g. in the case of the example above, if agent 3 had their turn in between agent 1 and 2, agent 3 would be able to initiate, or subscribe to, a different convention (since they do not meet any conditions of $c_{1}$) and $c_{1}$ will remain active. When more conventions are introduced, the learning problem becomes apparent and crucial, since agents must learn which convention is applicable in certain scenarios where more than one convention is available. Furthermore, since agents have the ability to opt in to a convention, they can also choose not to, which could result in an active convention termination without a completing action. This allows agents to halt certain conventions based on their unique observations alluding to it being non-optimal. Thus, the agents must learn when to prioritise superior conventions over existing or active inferior ones.
4.1 Architecture Design and Implementation
Algorithm 1 can be adapted to include conventions, and produce Algorithm 2, which serves as the pseudocode for conventions applied to independent Q-learning in a turn-based scenario. The current observation, which also contains the environment actions of the last round, determines the available and/or active conventions, shown in lines 9-15. Once all the conventions are determined, agent $i$ can choose to start or continue a convention based on the policy derived from $Q^{i}$ and their observation $O^{i}_{t}$, shown in line 16. Similar to human-play, the agents aren’t allowed to communicate which convention is currently being started or followed, but rather only observes the environment action determined by the appropriate convention policy, shown in line 22.
Algorithm 2 Independent Q-learning with conventions in a turn-based environment
1:Initialise hyperparameters similar to Algorithm 1
2:Define the list of conventions $\boldsymbol{C}$ with $C=\langle I,\pi_{I},F,\pi_{F}\rangle$
3:Initialise all $Q(O,C)\leftarrow 0$
4:for each episode do
5:     Reset environment and set total time $t\leftarrow 0$
6:     while $S_{t+1}$ is not terminal do
7:         for all players $i$ from 00 to $P-1$ do
8:              Obtain $O^{i}_{t}$ from environment
9:              for each convention $c$ in $\boldsymbol{C}$ do
10:                  if $O^{i}_{t}\in I_{c}$ then
11:                       Convention $c$ is available
12:                  else if $O^{i}_{t}\in F_{c}$ then
13:                       Convention $c$ is active
14:                  end if
15:              end for
16:              $C^{i}_{t}=\text{argmax}_{c}Q^{i}(O^{i}_{t},c)$triangleright Choose $C^{i}_{t}$ from the available and/or active conventions $\boldsymbol{c}$ based on $O^{i}_{t}$ and the policy derived from $Q^{i}$
17:              if $C^{i}_{t}$ was available then
18:                  $A^{i}_{t}=\pi_{Ic}(O^{i}_{t})$triangleright $\pi_{Ic}$ determines the environment action $A^{i}_{t}$
19:              else if $C^{i}_{t}$ was active then
20:                  $A^{i}_{t}=\pi_{Fc}(O^{i}_{t})$triangleright $\pi_{Fc}$ determines the environment action $A^{i}_{t}$
21:              end if
22:              Take action $A^{i}_{t}$ in the environment
23:              Receive and store $R_{t+1}$ and $O^{i}_{t+1}$, with $O^{i}_{t+1}\in S_{t+1}$, from environment
24:              $\sum_{k=0}^{P-1}R_{t+1+k}$triangleright Calculate the FAR based on the rewards for the next round of actions resulting from $A^{i}_{t}$
25:              $Q^{i}(O^{i}_{t},C^{i}_{t})\leftarrow Q^{i}(O^{i}_{t},C^{i}_{t})+\alpha\bigr{[}\sum_{k=0}^{P-1}R_{t+1+k}+\gamma\max_{c}Q^{i}(O^{i}_{t+P},c)-Q^{i}(O^{i}_{t},C^{i}_{t})\bigr{]}$
26:              $t\leftarrow t+1$
27:         end for
28:     end while
29:end for
From the environment’s perspective agents are merely choosing actions based on the observations the environment provided, but in reality the agents are choosing these actions based on predefined convention policies $\pi_{I}$ or $\pi_{F}$. In the case of deep RL, lines 9–14 and 17–21 indicate that our approach requires two new layers to be added to the agent’s architecture, namely one before the input layer to determine the available and/or active conventions and one at the output to translate the chosen convention into an environment action. This is illustrated in Fig. 2, and in the case of deep Q-learning, the action selection step is performed using $\text{argmax}_{c}Q(O_{t},c)$, with the available and/or active conventions acting as an action mask.
Figure 2: Architecture design for a feed forward neural network (NN) with conventions applied. The environment provides the observation $O_{t}$ and receives the action $A_{t}$. The size of the input layer is equal to the size of the observation tuple ($x$), and each hidden layer has a size equal to the number of atoms for that layer. The size of the output layer is equal to twice the number of conventions ($2y$), since we focus on two-step conventions, and the convention selection is determined by the MARL algorithm.
4.2 Action Space Augmentation
However, there can exist scenarios where no conventions are available, and the agents will not be able to take any actions. To solve this problem, we propose augmenting the action space with primitive actions and conventions simultaneously, similar to the discussion by Sutton et al. [19] on combining primitive actions and options. This requires that the conventions be appended to the existing primitive-action space (with size $z$), and results in an action-convention space consisting of
$C^{\prime}=\{a_{0},a_{1},\dots,a_{z},c_{1},c_{2},\dots,c_{y}\}.$ (6)
We can adapt Algorithm 2 to include an augmented action-convention space, and produce Algorithm 3.
Algorithm 3 Independent Q-learning with an augmented action-convention space in a turn-based environment
1:Initialise hyperparameters similar to Algorithm 1
2:Define the list of convention $\boldsymbol{C}$ with $C=\langle I,\pi_{I},F,\pi_{F}\rangle$
3:Combine primitive actions and conventions to produce the augmented action-convention space $C^{\prime}$
4:Initialise all $Q(O,C^{\prime})\leftarrow 0$
5:for each episode do
6:     Reset environment and set total time $t\leftarrow 0$
7:     while $S_{t+1}$ is not terminal do
8:         for all players $i$ from 00 to $P-1$ do
9:              Obtain $O^{i}_{t}$ from environment
10:              for each convention $c$ in $\boldsymbol{C}$ do
11:                  if $O^{i}_{t}\in I_{c}$ then
12:                       Convention $c$ is available
13:                  else if $O^{i}_{t}\in F_{c}$ then
14:                       Convention $c$ is active
15:                  end if
16:              end for
17:              ${C^{i}_{t}}^{\prime}=\text{argmax}_{ac}Q^{i}(O^{i}_{t},ac)$triangleright Choose ${C^{i}_{t}}^{\prime}$ from the available and/or active conventions $\boldsymbol{c}$ or primitive actions $A$ based on $O^{i}_{t}$ and the policy derived from $Q^{i}$
18:              if ${C^{i}_{t}}^{\prime}\in\boldsymbol{C}$ then $\triangleright$ If the chosen action is a convention
19:                  if $C^{i}_{t}$ was available then
20:                       $A^{i}_{t}=\pi_{Ic}(O^{i}_{t})$triangleright $\pi_{Ic}$ determines the environment action $A^{i}_{t}$
21:                  else if $C^{i}_{t}$ was active then
22:                       $A^{i}_{t}=\pi_{Fc}(O^{i}_{t})$triangleright $\pi_{Fc}$ determines the environment action $A^{i}_{t}$
23:                  end if
24:              else if ${C^{i}_{t}}^{\prime}\in A$ then $\triangleright$ If the chosen action is a primitive action
25:                  $A^{i}_{t}={C^{i}_{t}}^{\prime}$
26:              end if
27:              Take action $A^{i}_{t}$ in the environment
28:              Receive and store $R_{t+1}$ and $O^{i}_{t+1}$, with $O^{i}_{t+1}\in S_{t+1}$, from environment
29:              $\sum_{k=0}^{P-1}R_{t+1+k}$triangleright Calculate the FAR based on the rewards for the next round of actions resulting from $A^{i}_{t}$
30:              $Q^{i}(O^{i}_{t},C^{i}_{t})\leftarrow Q^{i}(O^{i}_{t},C^{i}_{t})+\alpha\bigr{[}\sum_{k=0}^{P-1}R_{t+1+k}+\gamma\max_{ac}Q^{i}(O^{i}_{t+P},ac)-Q^{i}(O^{i}_{t},C^{i}_{t})\bigr{]}$
31:              $t\leftarrow t+1$
32:         end for
33:     end while
34:end for
We can also adapt Fig. 2 to account for an augmented action-convention space by simply extending the output layer to include the primitive actions along with the conventions, and account for the difference between primitive action selection over convention selection in the convention-action translation. To explore and validate the capabilities of conventions and action space augmentation, we conduct preliminary tests using a simplified version of Hanabi, called Small Hanabi.
4.3 Preliminary Results on Small Hanabi
Small Hanabi has a deck reduced to 20 cards, i.e. only having two suits, is limited to only two players, and the player hand size has been reduced to two cards. Furthermore, the life tokens have been reduced to one, and the hint tokens to three, otherwise the game shares all the core mechanics as regular Hanabi. This results in the problem having a significantly smaller state-action space and allows for initial testing and development of algorithms, with the aim of solving the full Hanabi problem. Even though this problem is considered simpler than Hanabi, it is by no means an easy problem to solve, with a perfect score of 10/10 being difficult to achieve. Player hands are often “locked” with important cards that must not be discarded (called critical cards), hint tokens run out almost immediately, and players are often forced into no-win scenarios where they must choose a bad action or risk losing their single life token.
In these tests, we use our in-house learning environment as well as the open-sourced learning environment developed by Bard et al. [36]. We implement our own rudimentary conventions specific to Small Hanabi, which we developed through multiple human plays. Fig. 3(a) shows the result for independent Deep Q-learning (DQN) with a primitive-action space compared to independent Deep Q-learning with a pure convention [^10] A pure convention space accounts for all possible scenarios, and the agents will always have a convention available. In Small Hanabi this is possible due to the small state-action space, but becomes a far greater challenge in the full Hanabi problem. space. DQN with a pure convention space performs significantly better than DQN with a primitive-action space, achieving a faster training time and improved converged performance. To validate these results, and improve on the performance further, we conduct an initial test using Rainbow and the open-sourced Hanabi learning environment with the Small Hanabi preset, the results are shown in Fig. 3(b). Rainbow is able to perform better than independent Deep Q-learning achieving a score of 7,6/10 compared to 5/10, and when substituting the primitive-action space for a pure convention space, the agents achieve significantly faster training time along with a slight convergent performance increase.
Figure 3: (a) Learning curves for independent Deep Q-learning (DQN) with a primitive-action space, pure conventions space, and an augmented and simplified primitive-action-convention space tested in our in-house Small Hanabi environment. (b) Learning curves for Rainbow with a primitive-action space, pure conventions space, and an augmented and simplified primitive-action-convention space tested in DeepMind’s Hanabi learning environment with the Small Hanabi preset [36].
Due to Hanabi having a long list of intricate conventions, in addition to augmenting the action space with conventions, we experiment with the idea of simplifying the list of conventions during augmentation. This simplified list consists of the conventions that were the most straightforward and natural to develop, and were also the most common ones to occur in a game. The removed conventions can be considered edge-case or specific to certain game states that are very uncommon, and often results in uncertainty regarding their effectiveness. Fig. 3(a) and Fig. 3(b) shows the performance for DQN agents and Rainbow agents, respectively, with an augmented action-conventions-space. The performance impact is almost insignificant when compared to each method’s results for a pure convention space. This proves that even though only half of the conventions are present alongside primitive actions, the agents are still able to perform significantly better compared to only having primitive actions, and only slightly impacts performance compared to pure conventions. Therefore, when applying conventions in the full Hanabi problem, only a subset of established human conventions will be needed alongside primitive actions. Most existing Hanabi conventions have a list of core conventions which are built on the most essential principles, and are generally considered standardised. It is these conventions [^11] For an in depth discussion on each convention see the H-Group section I – VIII. alongside primitive actions that will be used when testing agents in the full Hanabi problem.
It is important to note that there could exist a combination of conventions that would result in better performance uplifts, especially when a larger variety of conventions are implemented. However for the sake of this argument, we focused on the principle conventions since they are generally the least contentious and offer a good starting point for exploring the capabilities of action space augmentation with conventions in Hanabi. Further research can be conducted to explore more advance conventions, including multistep conventions, and their potential performance benefits.
5 Performance Evaluation Using Hanabi
The preliminary tests for action augmentation with conventions in Small Hanabi has shown promising results, and allow the agents to achieve faster training time and improved policies. We now shift focus to the full Hanabi problem for 2–5 players and with the principle conventions defined by the H-Group. Our results will focus on Bard et al.’s [36] Rainbow agent as a baseline, and apply an augmented action-convention space to improve on its performance. Rainbow was chosen as our baseline since it is among the methods that train the fastest, is the most sample efficient among the existing Hanabi agents [36, 13, 14, 40, 27], has little run to run variance [36], is considered a widely applicable algorithm [57, 58, 59], and still leaves considerable performance to be desired within Hanabi. All tests are conducted on the open-sourced Hanabi learning environment developed by Bard et al. [36].
5.1 Experimental Setup
Before discussing the results, we will briefly highlight the architecture design for each deep RL technique. For a full list of the hyperparameters used in each method see Table 4 in Appendix A. All methods receive a one-hot encoded observation defined by the environment as input to their neural networks, notably this observation tuple also contains the most recent actions within the previous round. The Rainbow agents use a Multilayer Perceptron (MLP) consisting of two feed forward neural networks with 512 neurons each, with the ReLU activation function [60] applied. It implements the Adam optimizer [54] to calculate the TD error and perform backpropagations to update the network weights. Additionally, the Rainbow agents use distributional reinforcement learning to predict the value distributions which are approximated as a discrete distribution over 51 uniform atoms, along with prioritise replay memory sampling [28]. Even though Rainbow uses n-step bootstrapping, in these experiments a value of n=1 was found to be optimal, additionally noisy nets have been disabled in favour for a traditional decaying epsilon-greedy approach.
When applying conventions to Rainbow, we only augment the action space and leave the reward signal and observation space untouched. Additionally, we add the necessary convention layers to the network architecture, but keep the core algorithm unaltered, as illustrated by Fig. 2. Through testing, we found the optimal hyperparameters to be similar to baseline Rainbow, as shown in Table 4 in Appendix A. During evaluation we compare each approach’s exponential moving averages (with a weight value of 0.9995) and standard error of the mean. All agents are trained using an Intel Core i7 10 700K CPU and Nvidia RTX 3080Ti GPU. The reward type for the Hanabi learning environment is set to non-lenient, i.e. the agents will receive a large negative reward (equal to the score) when bombing out.
Due to the symmetric nature of the environment, we apply a shared policy strategy during training, i.e. the agents share an action-value function that is updated based on a communal memory of each agent’s experiences [61]. It is important to note that this still restricts learning by only using individual experiences, i.e., there is no additional sharing of state information between each agent. During the cross-play evaluation of 2–5 player Hanabi, we chose samples from 10 separately trained agents for each player count, and show the results for the combination of agents that performed the best on average.
5.2 Self-play Performance
The learning curves for Rainbow with and without an augmented action-convention space for 2–5 player Hanabi are shown in Fig. 4. The two player scenario seen in Fig. 4(a) shows the smallest improvement when applying conventions, however the two player Hanabi problem can be considered a special case. This is due to the fact that the shared information between the two players are far more limited when compared to higher player counts.
(a) 2 Player learning curves
(b) 3 Player learning curves
(c) 4 Player learning curves
(d) 5 Player learning curves
Figure 4: Learning curves with exponential moving averages (weight=0.9995) for Rainbow, obtained from Bard et al. [36], as baseline compared to Rainbow with an augmented action-convention space for Hanabi two to five players. The best agent is highlighted in black for each agent scenario within each player count.
For example, in two player Hanabi, half of the hands are hidden to a single player whereas in the three player scenario only a third of the hands are hidden, and the players always have a common hand visible. This allows for more advanced reasoning over another player’s actions, and can be seen in two of the principle conventions (specifically the Finesse and the Prompt) which cannot be applied in two player Hanabi.
Despite the fact that the complexity of the problem increases exponentially as more agents are added to the environment, the performance uplift of conventions become more apparent. In the three player setting seen in Fig. 4(b), the agents are able to achieve a significantly faster training time as well as an improved convergent performance with a higher average score. The performance improvement becomes even more apparent when looking at the five player setting, seen in Fig. 4(d), which is generally considered the most difficult problem to solve, and where the agents train roughly 10 times faster than the baseline Rainbow agents. In Fig. 5 the distribution of scores during evaluation of each agent over the course of 1 000 episodes is shown. In each scenario, conventions allow for a significant performance improvement, achieving a consistent grouping near a score of 21/25 and less overall variance when compared to baseline Rainbow. Rainbow with an augmented action-convention space also displays a robustness to increased player counts, with a remarkably consistent behaviour across all scenarios.
(a) 2 Player
(b) 3 Player
(c) 4 Player
(d) 5 Player
Figure 5: Distribution of scores for 2–5 player Hanabi over the course of 1 000 evaluation episodes comparing baseline Rainbow and Rainbow with an augmented action-convention space. The baseline Rainbow results were obtained from Bard et al. [36].
Table 1 shows the average score for the best agent from each player count’s various runs compared to the results from Bard et al. [36]. Rainbow with an augmented action-convention space demonstrates an improved score for each scenario, with a higher average in the 3–5 player scenario and overall less deviation. Additionally, the four and five player results are competitive compared to other algorithms, such as ACHA, which requires 20 billion training steps to achieve a score of 16.8/25 in the five player setting.
Table 1: Performance evaluation over 1 000 episodes for the best performing Rainbow agent with an augmented action-convention space applied taken from a sample of 10 different runs for each player count, compared to the results presented by Bard et al. [36]. Each score is averaged out of 25 with the standard error of the mean shown in brackets, and the number of perfect games (25/25) shown as a percentage.
Method 2P 3P 4P 5P
Rainbow with Conventions 20.65 (0.11) 3.6% 20.32 (0.07) 0.2% 20.09 (0.09) 0.1% 19.05 (0.08) 0%
Rainbow [36] 20.64 (0.22) 2.5% 18.71 (0.20) 0.2% 18.00 (0.17) 0% 15.26 (0.18) 0%
ACHA [36] 22.73 (0.12) 15.1% 20.24 (0.15) 1.1% 21.57 (0.12) 2.4% 16.80 (0.13) 0%
Upon investigation, the agents tend to choose conventions over primitive actions 70% of the time. This is the main reason for the Rainbow agents not achieving a higher average score, since they struggle to learn some of the core conventions available to them (specifically the Chop). We believe this to be a result of the Rainbow algorithm not having a memory of past observations, and therefore not being able to realise the importance of conventions that have a high payoff over extended time frames.
5.3 Cross-play Performance
Although self-play has historically led to the highest state-of-the-art agent performance within Hanabi [14], cross-play is also a crucial and widely applicable problem to solve. The ability for agents to cooperate across different training runs or regimes, or even across architectures, can greatly benefit MARL algorithms and their ability to cooperate in real world scenarios [47, 48]. Self-play agents are notorious for not being able to cooperate with never-before-seen partners, and often results in very poor performance [36, 27, 47, 48].
Table 2 shows the two player performance for various agents in cross-play Hanabi taken from existing literature, compared to the results of our cross-play Rainbow agents with an augmented action-convention space. Conventions are able to significantly improve on the performance of cross-play agents, allowing the baseline Rainbow agents to go from abysmal performance to an acceptable score. Furthermore, it is able to outperform SAD with only other-play applied for the two player scenario, with other-play and auxiliary task applied to SAD still holding the best score. However, as mentioned previously, auxiliary tasks only benefit the two player scenario in Hanabi [13]. When moving from self-play to cross-play, Rainbow with an augmented action-convention space only suffered a slight performance decrease and maintained remarkable consistency across various agent parings.
Table 2: Performance evaluation in two player Hanabi for the best performing agents in a cross-play scenario over the course of 1 000 episodes. The Rainbow with an augmented action-convention space agents were chosen from the top performing two agents in 10 different runs. Each score is averaged out of 25 with the standard error of the mean shown in brackets.
Method 2P
Self-play (SP) Cross-play (CP)
Rainbow with Conventions 20.65 (0.11) 17.02 (0.25)
Rainbow [36] 20.64 (0.22) 2.91 (1.67)
ACHA [36] 22.73 (0.12) 3.31 (1.78)
SAD [27] 23.94 (0.04) 2.52 (0.34)
SAD with other-play [27] 23.93 (0.02) 15.32 (0.65)
SAD with auxiliary tasks and other-play [27] 24.06 (0.02) 22.07 (0.11)
Table 3 shows the results for cross-play Rainbow with an augmented action-convention space for all player counts of Hanabi. As seen in the self-play results for Rainbow conventions and confirmed in Table 3, the two player scenario is a special case, with the agents showing the biggest performance degradation when moving from self-play to cross-play, and overall worst score compared to the other cross-play results. This is again due to the more powerful conventions not being available in two player Hanabi, thus, it is important to look at all the player scenarios when evaluating cross-play performance.
Table 3: Performance evaluation in 2–5 players Hanabi for the best performing Rainbow with an augmented action-convention space agents in a cross-play scenario over the course of 1 000 episodes. Each combination of agents in each player count were chosen from the top performing agents in a sample of 10 different runs. Each score is averaged out of 25 with the standard error of the mean shown in brackets.
Method 2P 3P 4P 5P
SP CP SP CP SP CP SP CP
Rainbow with Conventions 20.65 (0.11) 17.02 (0.25) 20.32 (0.07) 18.60 (0.15) 20.09 (0.09) 18.56 (0.11) 19.05 (0.08) 17.69 (0.09)
Most remarkable is the five player results, where the cross-play Rainbow agents with an augmented action-convention space are able to achieve a relatively high score and still cooperate effectively. The reason conventions are able to offer such a large performance gain for cross-play is due to the fact that the agents learn from a communal list of conventions that are taken from existing domain knowledge. Thus, if an agent is paired with a never-before-seen cooperator that has learned to play with the same list of conventions, they will still be able to cooperate effectively, since there is no uncertainty over another agent’s intent and reasoning. This is not unexpected, since humans who learn from the same “list” of conventions (whether it is physical or passed down), and that have never before interacted, can cooperate and communicate effectively with relatively low uncertainty.
6 Conclusion
Hanabi offers a complex and intricate problem for multi-agent reinforcement learning agents. It tests the ability to reason over another player’s actions and intensions, while maintaining a limited communication channel and partial observability. Existing MARL algorithms focus on complex architecture design and implement advanced algorithms capable of achieving remarkable performance. However, these algorithms are often computationally expensive, and require immense amounts of training data to learn convergent policies. In this paper, we focused on a different approach, exploring another core aspect of the Hanabi problem, namely conventions, and how to incorporate them into MARL. We showed how existing human conventions can be implemented in a MARL scenario using a special form of cooperative actions and action space augmentations.
Our main results show that conventions are able to significantly improve on the training time for Rainbow agents in Hanabi, and in the case of three to five players, also improve on the converged policy. Other Hanabi algorithms, such as SAD or MAPPO, require billions of training steps to reach convergent policies, whereas conventions reduced the number of training steps for Rainbow to below 30 million consistently. Additionally, conventions showed significant performance increase for cross-play agents, and allowed the Rainbow agents to go from not cooperating at all to being able to achieve decent performance across all player scenarios. For the most difficult scenario of five players, the cross-play Rainbow conventions agents trained on 50 million steps are able to outperform the ACHA self-play agents trained on 20 billion steps.
The biggest avenue for future research is exploring and developing convention discovery, similar to that of option discovery, which will allow agents to produce and define their own conventions as they train and learn. Our research serves as the basis for this future endeavour, proving the importance of conventions within MARL problems containing partial observability and limited communication.
[…]
Acknowledgements
This research was funded in part by the DW Ackerman Bursary as distributed by Stellenbosch University.
[…]
Author contributions
All authors contributed to the study conception and design. Material preparation and data collection were performed by F. Bredell, and all authors contributed equally to analysing the data. The first draft of the manuscript was written by F. Bredell and all authors commented on previous versions of the manuscript. All authors read and approved the final manuscript.
Declarations
Appendix A Hyperparameters
Herein follows a list of the hyperparameters used for independent deep Q-learning, Rainbow, and their conventions variants shown in Table 4. Each method’s hyperparameters were optimized using parameter sweeps, i.e., training each method with various combinations of hyperparameters over the course of multiple runs and selecting the hyperparameters with the best results.
Table 4: Table of hyperparameter values for each method used in our evaluation of Small Hanabi and Full Hanabi 2–5 Players. The hyperparameters used for Rainbow were sourced from Bard et al. [36], and after testing found to be the same after implementing conventions.
Deep Q-learning (With and without conventions)
Hyperparameter Value Description
Learning rate $\alpha$ 0.00005 Learning rate used by Adam optimizer
Discount factor $\gamma$ 0.5 Discount factor used in Q-learning update step
Training exploration $\epsilon_{t}$ 0.001 Value of $\epsilon$ in $\epsilon$-greedy strategy during training
Evaluation exploration $\epsilon_{e}$ 0.00 Value of $\epsilon$ in $\epsilon$-greedy strategy during evaluation
Training exploration decay 1 000 Decay rate of $\epsilon$ starting at 1 and ending in $\epsilon_{t}$
Replay memory size 35 000 Where experiences are stored to be used in the update step
Sampling batch size 64 Number of experiences randomly sampled from the experience replay memory used in the update step
Target network update frequency 200 Number of steps before the target network is updated with the policy network
Rainbow (With and without conventions)
Hyperparameter Value Description
Learning rate $\alpha$ 0.000025 Learning rate used by Adam optimizer
Discount factor $\gamma$ 0.99 Discount factor used in Q-learning update step
Training exploration $\epsilon_{t}$ 0.00 Value of $\epsilon$ in $\epsilon$-greedy strategy during training
Evaluation exploration $\epsilon_{e}$ 0.00 Value of $\epsilon$ in $\epsilon$-greedy strategy during evaluation
Training exploration decay 1 000 Decay rate of $\epsilon$ starting at 1 and ending in $\epsilon_{t}$
Replay memory size 50 000 Where experiences are stored to be used in the update step
Sampling batch size 32 Number of experiences randomly sampled from the experience replay memory used in the update step
Target network update frequency 500 Number of steps before the target network is updated with the policy network
Distribution atoms 51 Number of distributional atoms over which the value distributions are approximated as a discrete distributions
Step count $n$ 1 Number of steps over which the return is constructed in n-step bootstrapping